OpenStudy (anonymous):

I will reward with fan and medal for whoever helps. Which number would you add to three and nine-tenths to give an approximate sum of 6? Two and one-thirds Three and two-elevenths One and one-twelfths Two and seven-eighths

OpenStudy (anonymous):

\[6 = 3\frac{ 9 }{ 10 } + x\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you set your fraction back into and improper and subtract it from both sides, \[3\frac{ 9 }{ 10 } = \frac{ 39 }{ 10 }\] \[6 = \frac{ 39 }{ 10 } + x\] \[6 - \frac{ 39 }{ 10 } = x\]

OpenStudy (anonymous):

Now finding the least common denominator to be 10 we set 6 as an improper fraction. \[6 = \frac{ 60 }{ 10}\] So that we now have: \[\frac{ 60 }{ 10 } - \frac{ 39 }{ 10 } = x\] If we subtract we find the x. Where x is the missing number we would add to get 6

OpenStudy (anonymous):

I still dont understand that last part

OpenStudy (anonymous):

Which last part? I changed the 6 into an improper fraction so that we would have a common denominator so that they will be able to subtract from one another.

OpenStudy (anonymous):

So if you subtract: \[\frac{ 60 }{ 10 } - \frac{ 39 }{ 10 }\] What do we have?

OpenStudy (anonymous):

21/10

OpenStudy (anonymous):

Correct and we can simplify that further because we know that 10 goes into 21 2 times.

OpenStudy (anonymous):

okay so the 2 will be the whole number in the mixed number?

OpenStudy (anonymous):

Yes, and then we will only have a 1 left so that we are left with: \[2\frac{ 1 }{ 10 }\]

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

No problem